5. The heats of combustion of methane and butane are 890 kJ/mol, and 2876 kJmol. respectively. How much heat will be produced by combusting 125 g of each? Which will generate more heat for the same volume of gas? Explairn.
Answer #1

The heat of combustion of 1 mol of methane = 890 kJ

1 mole of methane = 16 g of methane

Number of moles of methane in 125 g = 125/16 = 7.8125 moles

Heat of combustion of 125 g of methane = 7.8125*890 = 6953.125 kJ

The heat of combustion of 1 mol of butane = 2876 kJ/mol

1 mole of butane = 58.12 g

Number of moles of butane in 125g = 125/58.12 = 2.15 mol

Heat of combustion = 2.15*2876 = 6185.4 kJ

Butane would give more heat than methane for the given volume because it has more number of hydrogen atoms than methane.